# The Riemann Hypothesis is false

### v37

@@
where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [6, p.78]. Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (5) that
@@
$I_{1}=\frac{(T-H))(\Lambda(n))}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$, $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{\log(n/m)}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=-\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
@@
If $f(x)$ is a real function that is differentiable thrice and $f^{\prime}(x)$ is steadily decreasing on $[a,b]$, note that
$\sum_{a<n\leq b}g(n)e^{2\pi if(n)}\sim e^{-\frac{1}{4}\pi i}\sum_{\alpha<v\leq\beta}\frac{e^{2\pi i(f(x_{v})-vx_{v})}}{|f^{\prime\prime}(x_{v})|^{\frac{1}{2}}}g(x_{v})\ll\sum_{\alpha<v\leq\beta}\Big{|}\frac{g(x_{v})}{f^{\prime\prime}(x_{v})^{\frac{1}{2}}}\Big{|}$ (18)
where $f^{\prime}(x_{v})=v,\alpha=f^{\prime}(a)$ and $\beta=f^{\prime}(b)$ [6, p.78]. Taking $g(u)=1/u,f(u)=\frac{t\log(1+\frac{u}{m})}{2\pi},f^{\prime}(u)=\frac{t}{2\pi(m+u)},f^{\prime\prime}(u)=-\frac{t}{2\pi(m+u)^{2}},x_{v}=\frac{t}{2\pi v}-m,f^{\prime\prime}(x_{v})=\frac{v^{2}}{t}$ gives, for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$,
$\sum_{0<r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}=\sum_{0<r\leq\varepsilon m-1}\frac{e^{it\log(1+\frac{r}{m})}}{r}\ll\sum_{\alpha<v\leq\frac{t}{2\pi m(1+\varepsilon)}}\frac{\sqrt{t}}{\frac{t}{2\pi}-mv}\ll\sum_{\alpha<v\leq\log^{2}t}\frac{1}{\varepsilon\sqrt{t}}\ll T^{-1/4}\log^{2}T.$ (19)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll T^{3/4}\log^{2}T.$ (20)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll T^{3/4}\log^{2}T.$ (21)
@@
$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi n}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (22)

### v38

@@
where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78]. Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
@@
$I_{1}=\frac{(T-H))\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{\log(n/m)}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=-\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
@@
Note that if $f(t)$ is differentiable on $[a,b]$, $f^{\prime}(t)$ is monotonic decreasing on $[a,b]$, $f^{\prime}(a)=\beta,f^{\prime}(b)=\alpha,$ then
$\sum_{a<r\leq b}e^{2\pi if(r)}=\sum_{|v|\leq N}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x+o(1)$ (18)
as $N\rightarrow\infty$, where
$\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x\ll\Big{|}\frac{1}{\alpha-v}-\frac{1}{\beta-v}\Big{|}=\frac{\beta-\alpha}{|(\alpha-v)(\beta-v)|}$ (19)
see e.g. [6, p.10]. Taking $f(u)=\frac{t\log(1+\frac{u}{m})}{2\pi},a=0,b=m\varepsilon-1$ and $f^{\prime}(u)=\frac{t}{2\pi(m+u)}$ gives
$\sum_{-N\leq v\leq 0}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $\ll\sum_{-N\leq v\leq 0}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (20)
$=(\beta-\alpha)\sum_{0\leq v\leq N}\frac{1}{(\alpha+v)(\beta+v)}$ (21)
$\ll(\beta-\alpha)\sum_{0\leq v\leq N}\frac{1}{1+v^{2}}$ (22)
$\ll\beta-\alpha$ (23)
$\sim\frac{t\varepsilon}{2\pi m}.$ (24)
Choose a rational $t\in[T,H]$ so that $\alpha=\frac{t}{2\pi(m+M)}$ and $\beta=\frac{t}{2\pi m}$ are never integers, where $m\in\Big{[}\frac{T}{(1+\varepsilon)\log^{2}T},T(1+\varepsilon)\Big{]}$ is an integer and $M$ is the largest integer $\leq m\varepsilon-1$. Thus $v\neq\alpha$ in particular. Let $K_{0}$ be the largest integer $<\alpha$. We also have
$\sum_{1\leq v\leq\alpha}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ (25)
$\ll\sum_{1\leq v\leq K_{0}}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (26)
$\ll(\beta-\alpha)\int_{1}^{K_{0}}\frac{1}{(\alpha-v)(\beta-v)}\mathrm{d}v$ (27)
$\ll(\beta-\alpha)\log t$ (28)
$\ll(\beta-\alpha)\log t$ (29)
$\sim\frac{t\varepsilon}{2\pi m}\log t.$ (30)
Let $K_{1}$ be the least integer $>\beta$. Note that
$\sum_{\beta<v\leq N}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $\ll\sum_{K_{1}\leq v\leq N}\frac{\beta-\alpha}{(v-\alpha)(v-\beta)}$ (31)
$\ll(\beta-\alpha)\int_{K_{1}}^{N}\frac{1}{(v-\alpha)(v-\beta)}\mathrm{d}v$ (32)
$\ll(\beta-\alpha)\log t$ (33)
$\ll(\beta-\alpha)\log t$ (34)
$\sim\frac{t\varepsilon}{2\pi m}\log t.$ (35)
Since $\beta-\alpha\sim\frac{t\varepsilon}{2\pi m}\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$, note that the range $\alpha<v\leq\beta$ contains no integers. Thus
$\sum_{\alpha<v\leq\beta}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=0.$ (36)
Combining (36), (35), (30), (25) and (19), yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$ and by Abel summation it follows that
$\sum_{a<r\leq b}\frac{1}{r}e^{2\pi if(r)}=\sum_{0<r\leq\varepsilon m-1}\frac{1}{r}e^{it\log(1+\frac{r}{m})}=o(1).$ (37)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (38)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (39)
@@
$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (40)